2. Starting from one oasis, a camel walks 25 km in a direction 30° South of West and then walks 30 km toward the North to a second oasis. What distance separates the two oases? You may use triangle method or component method.


1
Expert's answer
2022-02-02T05:14:05-0500

The distance is

"D_{1N}=-25\sin30°=-12.5\text{ km},\\
D_{1W}=25\cos30°=21.7\text{ km},\\
D_{2N}=30\text{ km},\\
D_N=-12.5+30=17.5\text{ km},\\
D_W=21.7\text{ km},\\
D=\sqrt{D_W^2+D_N^2}=27.9\text{ km}."